OpenStudy (anonymous):

the graphic of a quadratic function y=f(x) has the maximum at point (3,4). if f(5)=-4, then f(x)=?

OpenStudy (amoodarya):

suppose f(x)=ax^2+bx+c a<0 now f(3)=4 f'(3)=0 f(5)=-4 put them and solve the system of equation

OpenStudy (anonymous):

i cant understand??

OpenStudy (amoodarya):

you know derivation ?

OpenStudy (amoodarya):

if not use this \[f(x)=a(x-b)^2 +c\]

OpenStudy (anonymous):

if i substitute x in the equation tere will be 3 unknowns a b and c??

OpenStudy (amoodarya):

you know derivation ?

OpenStudy (amoodarya):